# How Good is ChatGPT in Giving Advice on Your Visualization Design?

### v3

@@ 3.4.2. Further Observations in Human and ChatGPT Responses
—Most questions had a single response (84), while 35 questions had multiple responses. Average ratings generally favored responses with multiple contributors across all categories, except for topicality. Significant rating differences were observed in breadth, depth, and actionability. Single-respondent questions averaged 3.06 ( = 1.05) in breadth, 3.35 ( = 1.05) in depth, and 3.72 ( = 0.96) in actionability. Conversely, multi-respondent questions scored higher, averaging 4.06 ( = 0.98) in breadth, 3.94 ( = 0.98) in depth, and 4.26 ( = 0.74) in actionability. When there’s only one response to a question, the quality of that response becomes heavily reliant on the individual respondent, leading to significant variability in quality, as seen in Figure 3A (high quality) versus Figure 3C (low quality). However, when multiple respondents contribute, they complement each other and compensate for areas in which one respondent might fall short (Figure 3B).
@@
—For ChatGPT responses, the differences in ratings between feedback and visualization guideline questions consistently remained small (less than 0.2 rating points). However, Human responses received higher scores for addressing guideline questions and lower scores for addressing design feedback questions across all metrics. For instance, Human responses excelled in breadth and depth for visualization guideline questions, with average ratings of 4.09 ( = 0.95) for breadth and 4.25 ( = 0.97) for depth. In contrast, design feedback questions received lower average ratings, with 3.08 ( = 1.06) for breadth and 3.25 ( = 0.97) for depth.
Guideline questions frequently require broader perspectives and readily accessible design knowledge, with 84.4% (27/32) of their responses citing external references, in contrast to design feedback questions, where only 18.39% (16/87) of Human responses did so. Conversely, feedback questions entail an understanding of domain-specific data and tasks, potentially making it challenging to offer comprehensive insights. For instance, a question like “Is this color scheme suitable?” requires a deep understanding of the domain. Conversely, questions such as the one about gendered colors in visualizations555visguides:use-of-gendered-colours-in-visualization can easily elicit a variety of various viewpoints and existing resources.
@@
ChatGPT’s color vision deficiency is usually not problematic—Color-related questions were prominent. Out of the 119 questions, 32 were centered on color-based design feedback (e.g., “Is the choice of the colour scheme appropriate?” 777visguides:area-chart-generation-capacity-over-the-years). Users rarely mentioned the specific colors used in their visualizations, often providing only visual encoding information. As a result, ChatGPT’s responses were typically based on general color principles. Despite this limitation, ChatGPT surpassed its overall average in most categories, except for depth and actionability.
@@ 3.5. Takeaways
Our analysis reveals ChatGPT’s ability to respond to data visualization queries compared to Human counterparts. ChatGPT excelled in breadth, exploring a wide range of ideas and concepts, surpassing Human responses. Clarity was another area where ChatGPT scored highly, benefiting from its structured response format. However, ChatGPT’s suggestions often leaned towards being more generic. In contrast, Human experts often offered more tailored and specific advice; they enhanced their responses by providing reference resources and complementing each other in multi-respondent scenarios. While these findings are promising, they do not provide insights into how practitioners perceive the value of Human’s responses, prompting the need for further research.
@@ 4. Perception of practitioners toward ChatGPT’s utility
Building upon our previous research, we conducted a comparative interview study involving feedback sessions with ChatGPT and Human experts. We aimed to gain a deeper understanding of how data visualization practitioners perceive ChatGPT’s role as a design companion, uncovering their attitudes and perceptions regarding ChatGPT’s design feedback.
@@ 4.2. Participants
The study involved 12 participants (P10 to P21), creating a diverse group with varying professional backgrounds, experience levels, and familiarity with ChatGPT (see Table 1). Their roles encompassed a broad spectrum, including developers, managers, journalists, consultants, product designers, analysts, scientists, and students. Professional experience ranged from less than one year to over ten years in their respective fields. Furthermore, participants’ engagement with ChatGPT varied, with some being daily users while others interacted with it only occasionally or as the need arose.
@@ 4.3. Tasks & Procedures
Each interview began with a brief introduction by the moderator and was divided into three segments: a visualization feedback session with ChatGPT, a similar session with Human expert(s), and an open-ended interview. Six participants initiated the process with the ChatGPT session, while the other six commenced with the Human expert session. After each feedback session, we administered a survey. In five sessions, two visualization experts provided feedback, while in the remaining seven sessions, we had one expert present. Both experts are current professors in the field of visualization.
@@ 4.6. Takeaways
The findings from our study underscore the distinct strengths and limitations of ChatGPT in comparison to Human experts. The clear preference for Human experts among participants highlights their superior ability to provide tailored, context-specific advice and engage in more interactive and collaborative conversations. This preference is rooted in the Human experts’ deep understanding of visualization context and their capacity for nuanced, bespoke feedback.
@@ 5. Discussions
5.1. Synthesizing the results of first and second studies
Both studies consistently highlighted similar strengths and weaknesses of ChatGPT compared to Human counterparts. A notable difference was that participants generally showed a stronger preference for Human experts in the second study. This disparity might be attributed to several factors. One reason could be that while any member of the VisGuides community can respond to questions on the forum, the feedback sessions were consistently answered by experienced visualization experts. In generall, the responses from Human experts in these sessions closely paralleled the high-scoring Human responses observed in the Visguides. Moreover, human-to-human interaction might have also played a role. For instance, Human experts ensured that they did not overlook any aspects of participants’ questions, benefiting from real-time conversations. Overall, the second study might demonstrate that human perception of usefulness might extend beyond just response quality.
@@ 5.2. Limitations of our studies and future opportunities
While our studies were designed and conducted with considerable care, it is important to acknowledge certain constraints. One notable restriction is the lack of demographic and expertise data for the Human respondents in the VisGuides forum, which poses a challenge in making direct comparisons with the Human experts. Also, we relied on the experiential opinions of participants to assess the feedback sessions, providing more direct and accurate evaluations. Due to the dynamic nature of the feedback sessions, we were not able to apply the static evaluation metrics used in the VisGuides. Despite this limitation, our results were highly consistent.
In this work, we used a fixed version of ChatGPT, a fine-tuned iteration of GPT 3.5. As with many other new technologies, LLMs are continuously advancing. ChatGPT and other LLM-based chatbots are evolving to incorporate more recent data and support new features, including the ability to handle more nuanced language or to understand and generate images, although some of these features are only available to paid users. As a result, responses to the questions in our study will vary among newer LLMs: they can be different, better, or sometimes even worse in certain cases. However, our study results go beyond transient outcomes, as they provide a framework for assessing LLM responses and understanding human-LLM interactions.
Going forward, comprehensive benchmarking of LLMs in data visualization would be beneficial, akin to other domain-specific benchmarks (Fei et al., 2023; Jin et al., 2021; Zhang et al., 2023a). As we anticipate a surge in LLM-driven visualization advancements, ensuring these advancements are grounded in a profound understanding of LLM knowledge capacities would be helpful. Our collection of questions and evaluation metrics can be a stepping stone for future scalable benchmarking for various models beyond ChatGPT. LLM-based evaluation can further address tedious human evaluation issues (Zheng et al., 2023; Liu et al., 2023; Wang et al., 2023; Fu et al., 2023; Kim et al., 2023). This direction in assessing LLM capabilities in data visualization might expand to parallel the evaluation of human abilities in visual perception and cognition (Haehn et al., 2018; Yang et al., 2023).
5.3. Design considerations for LLM-based feedback systems
@@
5.4. ChatGPT for design knowledge education
@@ 6. Conclusion
In this study, we explored the potential of ChatGPT as a design companion, focusing on its capacity to provide valuable visualization insights to practitioners. Our findings reveal that ChatGPT capitalizes on a vast knowledge repository to generate diverse and imaginative suggestions. While limitations like the contextual depth of its responses and truthful interaction dynamics exist, participants expressed optimism about its future utility. For future work, we plan to investigate ways to build a design feedback system based on the lessons learned. This research direction will involve enhancing chart comprehension and enriching the knowledge base of LLMs with well-founded rationales, including visualization examples and empirical evidence, developing user interfaces and interaction designs tailored to the designer’s workflow, and evaluating the system’s practical applicability.
@@ Acknowledgements.
- Bartolomeo et al. (2023) Sara Di Bartolomeo, Giorgio Severi, Victor Schetinger, and Cody Dunne. 2023. Ask and You Shall Receive (a Graph Drawing): Testing ChatGPT’s Potential to Apply Graph Layout Algorithms. In EuroVis 2023 - Short Papers, Thomas Hoellt, Wolfgang Aigner, and Bei Wang (Eds.). The Eurographics Association. https://doi.org/10.2312/evs.20231047
@@
- Haehn et al. (2018) Daniel Haehn, James Tompkin, and Hanspeter Pfister. 2018. Evaluating ‘graphical perception’with CNNs. IEEE transactions on visualization and computer graphics 25, 1 (2018), 641–650.
@@
- Hutchinson et al. (2023) Maeve Hutchinson, Aidan Slingsby, Radu Jianu, and Pranava Madhyastha. 2023. Towards Visualisation Specifications from Multilingual Natural Language Queries using Large Language Models. (2023). https://doi.org/10.2312/evp.20231072
@@
- Jin et al. (2021) Di Jin, Eileen Pan, Nassim Oufattole, Wei-Hung Weng, Hanyi Fang, and Peter Szolovits. 2021. What disease does this patient have? a large-scale open domain question answering dataset from medical exams. Applied Sciences 11, 14 (2021), 6421.
@@
- Liu et al. (2023) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu. 2023. Gpteval: Nlg evaluation using gpt-4 with better human alignment. arXiv preprint arXiv:2303.16634 (2023).
@@
- Paik (2023) Soonk Paik. 2023. How I Created a Data Visualization With Zero Coding Skills, Thanks to ChatGPT. Nightingale magazine by the Data Visualization Society (4 April 2023). https://nightingaledvs.com/data-visualization-using-chatgpt-to-code/
@@
- Wang et al. (2023) Jiaan Wang, Yunlong Liang, Fandong Meng, Haoxiang Shi, Zhixu Li, Jinan Xu, Jianfeng Qu, and Jie Zhou. 2023. Is chatgpt a good nlg evaluator? a preliminary study. arXiv preprint arXiv:2303.04048 (2023).
@@
- Yang et al. (2023) Fumeng Yang, Yuxin Ma, Lane Harrison, James Tompkin, and David H. Laidlaw. 2023. How Can Deep Neural Networks Aid Visualization Perception Research? Three Studies on Correlation Judgments in Scatterplots. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 822, 17 pages. https://doi.org/10.1145/3544548.3581111

### v4

@@ 3.4.2. Further Observations in Human and ChatGPT Responses
—Most questions had a single response (84), while 35 questions had multiple responses. Average ratings generally favored responses with multiple contributors across all categories, except for topicality. Significant rating differences were observed in coverage, breadth, depth, and actionability. Single-respondent questions averaged 4.25 ( = 1.02) in coverage, 3.06 ( = 1.05) in breadth, 3.35 ( = 1.05) in depth, and 3.72 ( = 0.96) in actionability. Conversely, multi-respondent questions scored higher, averaging 4.66 (*, = 0.67) in coverage, 4.06 (**, = 0.98) in breadth, 3.94 (**, = 0.98) in depth, and 4.26 (*, = 0.74) in actionability. When there’s only one response to a question, the quality of that response becomes heavily reliant on the individual respondent, leading to significant variability in quality, as seen in Figure 3A (high quality) versus Figure 3C (low quality). However, when multiple respondents contribute, they complement each other and compensate for areas in which one respondent might fall short (Figure 3B).
@@
—For ChatGPT responses, the differences in ratings between feedback and visualization guideline questions consistently remained small. For ChatGPT 4, there was a significant difference in the depth metric (), with guideline questions receiving higher ratings (Mean = 4.64, = 0.48) compared to feedback questions (Mean = 4.35, = 0.57). For ChatGPT 3.5, the only significant difference was observed in the coverage metric (), where feedback questions scored higher (Mean = 4.97, = 0.24) compared to guideline questions (Mean = 4.79, = 0.48). Across other metrics, no significant differences were detected, highlighting consistent performance regardless of the question type.
However, we observed more significant differences in Human responses between guideline and design feedback questions. For the breadth metric, guideline questions scored significantly higher (Mean = 4.12, = 0.95) compared to feedback questions (Mean = 3.06, = 1.05, ). Similarly, for the coverage metric, guideline questions achieved a higher score (Mean = 4.76, = 0.43) than feedback questions (Mean = 4.22, = 1.05, ). The depth metric also showed a notable difference, with guideline questions receiving a substantially higher score (Mean = 4.24, = 0.95) compared to feedback questions (Mean = 3.24, = 0.98, ). Overall, Human responses to guideline questions tended to be more comprehensive and detailed. For other metrics, no significant differences were observed.
Guideline questions frequently require broader perspectives and readily accessible design knowledge, with 84.4% (27/32) of Human responses citing external references, in contrast to design feedback questions, where only 18.39% (16/87) of Human responses did so. Conversely, feedback questions entail an understanding of domain-specific data and tasks, potentially making it challenging to offer comprehensive insights. For instance, a question like “Is this color scheme suitable?” requires a deep understanding of the domain. Conversely, questions such as the one about gendered colors in visualizations555visguides:use-of-gendered-colours-in-visualization can easily elicit a variety of various viewpoints and existing resources.
@@
ChatGPT 3.5’s color vision deficiency was usually not problematic—Color-related questions were prominent. Out of the 119 questions, 32 were centered on color-based design feedback (e.g., “Is the choice of the colour scheme appropriate?” 777visguides:area-chart-generation-capacity-over-the-years). Users rarely mentioned the specific colors used in their visualizations, instead providing only visual encoding information in their questions. Consequently, ChatGPT 3.5 lacked access to exact color information due to its inability to comprehend images. Despite this limitation, there were no significant differences across all metrics in ChatGPT 3.5’s responses.
@@ 3.5. Takeaways
Our analysis shed light on ChatGPT’s cability to respond to data visualization queries compared to Human counterparts. ChatGPT clearly excelled in breadth, exploring a wide range of ideas and concepts, surpassing Human responses. Clarity was another area where ChatGPT scored highly, benefiting from its structured response format. ChatGPT 4 outperformed ChatGPT 3.5 by providing more focused and in-depth responses, aided by its image comprehension capabilities. Some Human responses resembled those of ChatGPT 4, often enhancing their quality by including reference resources and complementing one another in multi-respondent scenarios.
@@ 4. Perception of practitioners toward ChatGPT’s utility
While the VisGuides forum analysis findings are promising, they do not provide insights into how practitioners perceive the value of ChatGPT’s responses, prompting the need for further research. As a result, we conducted a comparative interview study involving feedback sessions with ChatGPT and Human experts. Experts were chosen over anonymous users to ensure the quality of human responses and establish a higher benchmark for evaluating the limitations of ChatGPT. These experts also represent experienced colleagues who provide design feedback in professional settings (Choi et al., 2023b; Esteves and Neves, 2022). Our goal was to explore how data visualization practitioners evaluate ChatGPT’s effectiveness as a design companion, focusing on their experiences and perspectives on its design feedback.
@@ 4.2. Participants
The study involved 12 participants (P10 to P21), creating a diverse group with varying professional backgrounds, experience levels, and familiarity with ChatGPT (see Table 2); we started participant codes from 10 to ensure that all IDs were consistently two digits long. Their roles encompassed a broad spectrum, including developers, managers, journalists, consultants, product designers, analysts, scientists, and students. Professional experience ranged from less than one year to over ten years in their respective fields. Furthermore, participants’ engagement with ChatGPT varied, with some being daily users while others interacted with it only occasionally or as the need arose.
@@ 4.3. Tasks & Procedures
Each interview began with a brief introduction by the moderator and was divided into three segments: a visualization feedback session with ChatGPT, a similar session with Human expert(s), and an open-ended interview. Our study employed a within-subjects design with counterbalancing to control for order effects. Six participants initiated the process with the ChatGPT session, while the other six commenced with the Human expert session. After each feedback session, we administered a survey. In five sessions, two visualization experts provided feedback, while in the remaining seven sessions, we had one expert present. Both experts are current professors in the field of visualization.
@@ 4.6. Takeaways
The findings from our study underscore the distinct strengths and limitations of ChatGPT in comparison to Human experts. The clear preference for Human experts among participants highlights their superior ability to provide tailored, context-specific advice and engage in more interactive and collaborative conversations. This preference is rooted in the Human experts’ deep understanding of visualization context and their capacity for nuanced, bespoke feedback. The feedback provided by Human experts in these sessions closely paralleled the high-scoring Human responses observed in the Visguides.
@@ 5.1. Limitations of our studies and future opportunities
While our studies were designed and conducted with considerable care, there exist certain limitations. First, the absence of demographic and expertise data for the Human respondents in the VisGuides forum makes direct comparisons with the Human experts challenging. However, the human baselines in both studies represent distinct real-world scenarios: receiving feedback from anonymous online platforms or from experienced colleagues in the workplace. Also, due to the dynamic nature of the feedback sessions in the second study, we were not able to apply the static evaluation metrics used in the VisGuides. Despite this limitation, our findings were highly consistent across the two studies.
We used a fixed version of ChatGPT. ChatGPT and other LLM-based chatbots are evolving to incorporate more recent data and support new features, including the ability to handle more nuanced language or to understand and generate images, although some of these features are only available to paid users. As a result, responses to the questions in our study can vary among newer LLMs. Likewise, the recently released advanced voice mode could help narrow the interaction modality gap identified in the second study, potentially improving the fluidity of conversations. That said, it still falls short of achieving human-like naturalness, such as exhibiting proactive questioning and avoiding passive behavior. However, our study results provide a valuable foundation for future research, offering both datasets, metrics, and methodologies for evaluating LLM responses in the context of visualization design feedback.
5.2. Towards automatic evaluation of LLM feedback capabilities
Facilitating ongoing monitoring of LLMs’ design feedback capabilities would be beneficial to ensure that both the development of LLMs and the systems built upon them remain well-grounded. The current reliance on human intervention to evaluate the quality of responses is inefficient. Recently, LLMs have demonstrated remarkable capabilities in evaluating natural language generation, showing greater alignment with human evaluations compared to traditional similarity-based or embedding-based metrics such as BLEU or BERTScore (Zheng et al., 2023; Liu et al., 2023b; Wang et al., 2023a; Fu et al., 2023; Kim et al., 2023).
| Metric | Response type | Mean diff | Gwet’s AC1 | Score dist | ||
| Depth | Human | -0.563 | 0.558 | 0.631 | 0.208 | |
| ChatGPT | 0.542 | 0.215 | 0.228 | 0.29 | ||
| Breadth | Human | -0.24 | 0.553 | 0.62 | 0.424 | |
| ChatGPT | -0.083 | 0.522 | 0.537 | 0.713 | ||
| Clarity | Human | -0.792 | 0.372 | 0.395 | 0.057 | |
| ChatGPT | -0.01 | 0.517 | 0.52 | 0.901 | ||
| Topicality | Human | -0.771 | 0.366 | 0.396 | 0.208 | |
| ChatGPT | 0.063 | -0.029 | -0.029 | 0.915 | ||
| Actionability | Human | -0.928 | 0.47 | 0.525 | 0.167 | |
| ChatGPT | 0.406 | 0.466 | 0.492 | 0.422 | ||
| Coverage | Human | -0.87 | 0.41 | 0.457 | 0.097 | |
| ChatGPT | 0.01 | 0.424 | 0.424 | 0.925 | ||
| All | All | -0.27 | 0.404 | 0.433 | 0.444 | |
| Human | -0.694 | 0.455 | 0.504 | 0.194 | ||
| ChatGPT | 0.154 | 0.353 | 0.362 | 0.694 |
We tested the viability of this idea by developing an evaluation prompt (see below) and used it to assess both Human and ChatGPT 3.5’s responses based on the six evaluation metrics outlined in Table 1.
In the experiment, we employed the in-context learning method (Brown et al., 2020; Li et al., 2024) using GPT-4o-mini via API calls in December 2024, leveraging the LLM’s few-shot learning capabilities to generalize evaluation tasks from the provided examples. Based on these settings, the evaluation task was conducted in the following steps: First, the Visguides dataset was split into a few-shot development set (20%) and a test set (80%). Second, for each test case and a metric, examples were selected from the few-shot development set, where we settled on after testing values within a range between 1 to 8 given the limit of context lengths in the GPT-4o-mini. The process for the example selection was based on two principles: 1) Score pooling for extracting diverse examples: To ensure a diverse pool of examples ranging from low to high quality, the examples in the few-shot development set were grouped into bins for unique scores from 1 to 5 based on their scores for the target metric. 2) Embedding similarity for retrieving relevant examples: Within each bin, the most similar examples based on embedding similarity based on Sentence Transformers (Reimers and Gurevych, 2019) were selected to help guide the model to choose the examples that are most semantically close to the input query, effectively “picking the best match” from the limited example set. The number of examples was determined proportionally to the size of the corresponding bin. The selected examples were incorporated into the prompt and provided to GPT-4o-mini, released in August 2024, as presented below:
You are an expert evaluator specializing in data visualization assessments.
Your task is to evaluate responses to visualization-related questions as closely
aligned with given examples, focusing specifically on the {metric} metric.
Here are specific guidelines for {metric}:
Before evaluating, carefully consider:
1. The specific visualization requirements in the question
2. How well the response addresses those requirements
3. The technical accuracy and appropriateness of the visualization approach
4. The practicality and implementability of the solution
{metric-specific evaluation rubric}
{metric-specific evaluation guideline}
Here are some example evaluations:
{few-shot examples}
Now evaluate this {response type} response based on the following metric:
Question: {question}
Response to evaluate: {response}
Evaluate the response on {metric} using a scale of 1-5 (1 being lowest, 5 being
highest). Your job is to make your evaluation as closely aligned as the given examples,
calibrating the score closely to the examples.
Metric definition:
- {metric}: {metric definition}
For the criterion, provide:
1. A numerical score (1-5)
2. A brief justification for the score
3. Specific examples from the response
Your response MUST be a valid JSON object with this exact structure:
"evaluation": {
"{metric}": {
"score": X,
"justification": "...",
"examples": "..."
In the prompt, the metric-specific information such as evaluation rubric and guideline (Table LABEL:tab:eval-rubrics and LABEL:tab:eval-guidelines), few-shot examples, metric definition (Table 1), and response type (Human or ChatGPT) are integrated for each test case accordingly.
We then compare the output scores to those provided by human evaluators and calculate the alignment using agreement measures and correlation coefficients. In our experiment, we use Kendall’s tau (), Spearman (), and Gwet’s AC1 coefficient as underlying measures. While Kendall’s tau was selected as one of the most widely used metrics, several studies (Hou, 2024; Liu et al., 2023a) have pointed out that this measure can lead to biased results when the distribution of the underlying scores is severely skewed, failing to capture the true alignment. This issue is particularly relevant in our case for metrics such as clarity, coverage, and topicality, especially for ChatGPT’s responses, where scores are skewed toward higher values (e.g., 4 or 5). To address this, Gwet’s AC1 (Gwet, 2001), which has been proposed and increasingly adopted to mitigate biased judgments, was used to complement Kendall’s tau.
Table 4 presents the outcomes of these measures for each evaluation metric. Overall, the results show that the LLM evaluator’s scores are fairly aligned with those of humans (, ) obtained across all metrics, given a well-known interpretation of Kendall’s tau and Spearman coefficient (Akoglu, 2018; Dancey, 2007) (0.4-0.6: moderate agreement). Specifically, the alignment was better captured for both Human and ChatGPT responses in the breadth (; ) and actionability metric (; ), and in the depth metric for Human responses (). For metrics with highly skewed score distributions, such as topicality, Kendall’s tau coefficients were relatively low for ChatGPT responses. In these cases, Gwet’s AC1 coefficient provided a more nuanced perspective, demonstrating a high alignment (). While our experiment shows promising capabilities of automatic evaluation using LLMs, it is worth to further investigate refining heuristics and developing prompts, for example, making prompts targeted to response types or providing clearer metric descriptions, in order to improve the alignment.
6. Design considerations for LLM-based feedback systems
@@
Toward a better design knowledge education
@@ 7. Conclusion
In this study, we explored the potential of ChatGPT as a design companion, focusing on its capacity to provide valuable visualization insights to practitioners. Our findings reveal that ChatGPT capitalizes on a vast knowledge repository to generate diverse and imaginative suggestions. While limitations like the contextual depth of its responses and truthful interaction dynamics exist, participants expressed optimism about its future utility.
For future work, we plan to investigate ways to build a design feedback system based on the lessons learned. This research direction will involve enhancing chart comprehension and enriching the knowledge base of LLMs with well-founded rationales, including visualization examples and empirical evidence, developing user interfaces and interaction designs tailored to the designer’s workflow, and evaluating the system’s practical applicability.
@@ Acknowledgements.
- Akoglu (2018) Haldun Akoglu. 2018. User’s guide to correlation coefficients. Turkish journal of emergency medicine 18, 3 (2018), 91–93.
- Alexander et al. (2024) Jason Alexander, Priyal Nanda, Kai-Cheng Yang, and Ali Sarvghad. 2024. Can GPT-4 Models Detect Misleading Visualizations?. In 2024 IEEE Visualization and Visual Analytics (VIS). IEEE, 106–110.
- Awasthi et al. (2023) Raghav Awasthi, Shreya Mishra, Dwarikanath Mahapatra, Ashish Khanna, Kamal Maheshwari, Jacek Cywinski, Frank Papay, and Piyush Mathur. 2023. HumanELY: Human evaluation of LLM yield, using a novel web-based evaluation tool. MedRXIV (2023), 2023–12.
@@
- Bendeck and Stasko (2024) Alexander Bendeck and John Stasko. 2024. An empirical evaluation of the gpt-4 multimodal language model on visualization literacy tasks. IEEE Transactions on Visualization and Computer Graphics (2024).
@@
- Chen et al. (2024) Nan Chen, Yuge Zhang, Jiahang Xu, Kan Ren, and Yuqing Yang. 2024. Viseval: A benchmark for data visualization in the era of large language models. IEEE Transactions on Visualization and Computer Graphics (2024).
- Choe et al. (2024) Kiroong Choe, Chaerin Lee, Soohyun Lee, Jiwon Song, Aeri Cho, Nam Wook Kim, and Jinwook Seo. 2024. Enhancing Data Literacy On-demand: LLMs as Guides for Novices in Chart Interpretation. IEEE Transactions on Visualization and Computer Graphics (2024).
@@
- Dancey (2007) CP Dancey. 2007. Statistics without maths for psychology. Prentice Hall.
@@
- Duan et al. (2024a) Peitong Duan, Chin-Yi Cheng, Gang Li, Bjoern Hartmann, and Yang Li. 2024a. UICrit: Enhancing Automated Design Evaluation with a UI Critique Dataset. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology. 1–17.
- Duan et al. (2024b) Peitong Duan, Jeremy Warner, Yang Li, and Bjoern Hartmann. 2024b. Generating Automatic Feedback on UI Mockups with Large Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–20.
@@
- Gu et al. (2024) Ken Gu, Madeleine Grunde-McLaughlin, Andrew McNutt, Jeffrey Heer, and Tim Althoff. 2024. How do data analysts respond to ai assistance? a wizard-of-oz study. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–22.
- Gwet (2001) Kilem Gwet. 2001. Handbook of inter-rater reliability. Gaithersburg, MD: STATAXIS Publishing Company (2001), 223–246.
@@
- Hou (2024) Xinmeng Hou. 2024. Mitigating Biases to Embrace Diversity: A Comprehensive Annotation Benchmark for Toxic Language. arXiv preprint arXiv:2410.13313 (2024).
@@
- Inala et al. (2024) Jeevana Priya Inala, Chenglong Wang, Steven Drucker, Gonzalo Ramos, Victor Dibia, Nathalie Riche, Dave Brown, Dan Marshall, and Jianfeng Gao. 2024. Data Analysis in the Era of Generative AI. arXiv preprint arXiv:2409.18475 (2024).
- Jiang et al. (2024) Che Jiang, Biqing Qi, Xiangyu Hong, Dayuan Fu, Yang Cheng, Fandong Meng, Mo Yu, Bowen Zhou, and Jie Zhou. 2024. On Large Language Models’ Hallucination with Regard to Known Facts. arXiv preprint arXiv:2403.20009 (2024).
@@
- Kim et al. (2024) Nam Wook Kim, Hyung-Kwon Ko, Grace Myers, and Benjamin Bach. 2024. ChatGPT in Data Visualization Education: A Student Perspective . In 2024 IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC). IEEE Computer Society, Los Alamitos, CA, USA, 109–120. https://doi.org/10.1109/VL/HCC60511.2024.00022
@@
- Li et al. (2024) Dawei Li, Bohan Jiang, Liangjie Huang, Alimohammad Beigi, Chengshuai Zhao, Zhen Tan, Amrita Bhattacharjee, Yuxuan Jiang, Canyu Chen, Tianhao Wu, et al. 2024. From Generation to Judgment: Opportunities and Challenges of LLM-as-a-judge. arXiv preprint arXiv:2411.16594 (2024).
@@
- Liu et al. (2023a) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu. 2023a. G-eval: Nlg evaluation using gpt-4 with better human alignment. arXiv preprint arXiv:2303.16634 (2023).
- Liu et al. (2023b) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu. 2023b. Gpteval: Nlg evaluation using gpt-4 with better human alignment. arXiv preprint arXiv:2303.16634 (2023).
- Lo and Qu (2024) Leo Yu-Ho Lo and Huamin Qu. 2024. How Good (Or Bad) Are LLMs at Detecting Misleading Visualizations? IEEE Transactions on Visualization and Computer Graphics (2024).
@@
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics. https://arxiv.org/abs/1908.10084
- Reiter (2018) Ehud Reiter. 2018. A structured review of the validity of BLEU. Computational Linguistics 44, 3 (2018), 393–401.
@@
- Scott and Moore (2006) Donia Scott and Johanna Moore. 2006. An NLG evaluation competition? eight reasons to be cautious. (2006).
- Shen et al. (2024) Shuyu Shen, Sirong Lu, Leixian Shen, Zhonghua Sheng, Nan Tang, and Yuyu Luo. 2024. Ask Humans or AI? Exploring Their Roles in Visualization Troubleshooting. arXiv preprint arXiv:2412.07673 (2024).
@@
- Sreedhar et al. (2024) Makesh Narsimhan Sreedhar, Traian Rebedea, Shaona Ghosh, Jiaqi Zeng, and Christopher Parisien. 2024. CantTalkAboutThis: Aligning Language Models to Stay on Topic in Dialogues. arXiv preprint arXiv:2404.03820 (2024).
@@
- Tian et al. (2024) Yuan Tian, Weiwei Cui, Dazhen Deng, Xinjing Yi, Yurun Yang, Haidong Zhang, and Yingcai Wu. 2024. Chartgpt: Leveraging llms to generate charts from abstract natural language. IEEE Transactions on Visualization and Computer Graphics (2024).
@@
- Vaithilingam et al. (2024) Priyan Vaithilingam, Elena L Glassman, Jeevana Priya Inala, and Chenglong Wang. 2024. DynaVis: Dynamically Synthesized UI Widgets for Visualization Editing. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–17.
@@
- Van der Lee et al. (2021) Chris Van der Lee, Albert Gatt, Emiel Van Miltenburg, and Emiel Krahmer. 2021. Human evaluation of automatically generated text: Current trends and best practice guidelines. Computer Speech & Language 67 (2021), 101151.
- Van Miltenburg et al. (2018) Emiel Van Miltenburg, Desmond Elliott, and Piek Vossen. 2018. Measuring the diversity of automatic image descriptions. In Proceedings of the 27th International Conference on Computational Linguistics. 1730–1741.
- Wang et al. (2024c) Chenglong Wang, Bongshin Lee, Steven Drucker, Dan Marshall, and Jianfeng Gao. 2024c. Data Formulator 2: Iteratively Creating Rich Visualizations with AI. arXiv preprint arXiv:2408.16119 (2024).
- Wang et al. (2023c) Chenglong Wang, John Thompson, and Bongshin Lee. 2023c. Data Formulator: AI-powered Concept-driven Visualization Authoring. IEEE Transactions on Visualization and Computer Graphics (2023).
- Wang et al. (2024b) H.W. Wang, J. Hoffswell, V.S. Bursztyn, and C.X. Bearfield. 2024b. How Aligned are Human Chart Takeaways and LLM Predictions? A Case Study on Bar Charts with Varying Layouts. IEEE Transactions on Visualization and Computer Graphics (2024).
- Wang et al. (2024a) Huichen Will Wang, Mitchell Gordon, Leilani Battle, and Jeffrey Heer. 2024a. DracoGPT: Extracting Visualization Design Preferences from Large Language Models. IEEE Transactions on Visualization and Computer Graphics (2024).
- Wang et al. (2023a) Jiaan Wang, Yunlong Liang, Fandong Meng, Haoxiang Shi, Zhixu Li, Jinan Xu, Jianfeng Qu, and Jie Zhou. 2023a. Is chatgpt a good nlg evaluator? a preliminary study. arXiv preprint arXiv:2303.04048 (2023).
- Wang et al. (2023b) Ke Wang, Houxing Ren, Aojun Zhou, Zimu Lu, Sichun Luo, Weikang Shi, Renrui Zhang, Linqi Song, Mingjie Zhan, and Hongsheng Li. 2023b. Mathcoder: Seamless code integration in llms for enhanced mathematical reasoning. arXiv preprint arXiv:2310.03731 (2023).